Adam is 12 years older than his sister Liz. Four years ago Adam was 4 times older than his sister. How old is each now?
Answer by Alan3354(69443)

Adam is 12 years older than his sister Liz. Four years ago Adam was 4 times older than his sister. How old is each now?
---------
A = L + 12
A-4 = 5*(L-4) = 5L-20
Sub for A
L+12 - 4 = 5L-20
L+8 = 5L-20
4L = 28
L = 7
A = 19